40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 23:32:07 2011, Aidan, Jamie, Configuration, LSC, AS PD and Camera installed AS_beam.jpgP1070546.JPG
    Reply  Fri Apr 15 22:57:38 2011, Aidan, Jamie, Configuration, LSC, AS PD and Camera installed 
Message ID: 4536     Entry time: Fri Apr 15 22:57:38 2011     In reply to: 4521
Author: Aidan, Jamie 
Type: Configuration 
Category: LSC 
Subject: AS PD and Camera installed 

AS port ITMX YAW  range where AS beam was visible = [-1.505, -1.225] - these extrema put the beam just outside of some aperture in the system -> set ITMX YAW to -1.365

ITMX PITCH range = [-0.7707, -0.9707] -> set to ITMX PITCH to -0.8707

Quote:

I spent some time tracking down the AS beam which had vanished from the AP table. Eventually, by dramatically mis-aligning SRM, PRM and ITMY, returning BS to its Jan 1st PITCH and YAW values and tweaking the ITMX alignment [actual values to follow], I was able to get an AS beam out onto the AP table. I verified that it was the prompt reflection off ITMX by watching it move as I changed the YAW of that optic and watching it stay stationary as I changed the YAW of ITMY.

Jamie and I then steered the beam through a 2" PLCX-50.8-360.6 lens and placed the RF PD (AS55) at the focus. Additionally, we installed the AS camera to observe the leakage field through a Y1S steering mirror (as shown in the attached diagram).

Currently the PD has power but the RF and DC outputs are not connected to anything at the moment.

Atm 2 by Steve

 

 

 

ELOG V3.1.3-